9. A sealed tank containing seawater to a height of 11.0 m also contains air above the water at a gauge pressure of 3.00 atm. Water flows out from the bottom through a small hole. How fast is this water moving?

1
Expert's answer
2022-06-13T08:24:06-0400

"p+\rho gh=\frac{\rho v^2}2,"

"v=\sqrt{\frac{2p}{\rho}+2gh}=28.3~\frac ms."